if 4-letter "words" are formed using the letters A, B, C,D,E, F,G, how many such words are possible for each of the following conditions:
a. no condition is imposed.
b. no letter can be repeated in a word.
c. each word must begin with the letter A and letters can be repeated.
d. the letter C must be at the end and letters can be repeated.
e. the second letter must be a vowel and letters can be repeated.

Found 2 solutions by sudhanshu_kmr, edjones:
Answer by sudhanshu_kmr(1152)


a)
no. of ways to form 4-letter words if no condition = 7 * 7 * 7 * 7 = 2401
because, at each place any one of 7 can be used.

b)
no. of ways to form 4-letter words without repetition = 7P4 = 840


c)
no. of ways to form 4-letter words begin with A = 7 * 7* 7 = 343
because, at each of remaining 3 places any one of 7 can be used.


d)
similarly as previous part, no. of ways = 343


e)
here 2 vowels, A and E.
no. of words when A is on second place = 343 (as previous part)
no. of words when E is on second place = 343
total no. of ways = 2 * 343 = 686




Sometimes there may be typing mistake in solution of a problem, so please ignore it. Understand the concept and try to solve the problem yourself. If there is problem related to concept, contact at
sudhanshu.cochin@yahoo.com or sudhanshu.cochin@gmail.com
Best of luck.......

Answer by edjones(8007)

a. 7^4=2,401
b. 7P4=840
c. 7^3=343
d. 343
e. 2*343=686
.
Ed